OpenStudy (anonymous):

4x+5y=13 5x+y-8=0 its substitution question

OpenStudy (anonymous):

\[5x+y-8=0\iff y=-5x+8\] substitute that in to the first equation, get \[4x+5(-5x+8)=13\] and solve for \(x\)

OpenStudy (anonymous):

thankyou ^_^

OpenStudy (blank ):

I got: x=9/7 y=11/7